Q amount of heat is supplied to one mole of and ideal gas at constant volume to increase its temperature by 1degree Celsius .if same amount of heat is supplied to the gas isobarically,t he increase in the internal energy of the gas molecule is....

Grade:12

1 Answers

Vikas TU
7 years ago
U = W + Q
U at constant volume = 3/2 * 1 * R * 1 = 3R/2 J
As W = 0 = 3R/2 –  Q
Q = 3R/2...........................................................................(1)
 
Now same amount of heat is supplied therfore,
U = Q + W
U = 3R/2 + W
where W = 1*R*1= R
U = 3R/2 + R => 5R/2 J